I dont understand whats a half life? How do i get this answer?
The amount of carbon 14 in a given organism at the time of that organism’s death (time zero) is 100%.
The half-life of carbon 14 is 5750 years, which means only 50% of the original carbon 14 will remain in
the given object 5750 years later. How old is an object that is found to contain only 20% of its original
carbon 14, to the nearest ten years?


Answer by rothauserc(4718)

carbon-14 is a radioactive isotope of carbon, half-life is the time for half the carbon-14 in an object to decay into nitrogen-14.
:
The appropriate formula is
:
A = Ao x (1/2)^(t / HL), A is the amount currently remaining, Ao is the original amount, t is the amount of time that has passed, and HL is the half-life
:
0.20 = 1.00 x (1/2)^(t / 5750)
:
0.20 = (1/2)^(t / 5750)
:
use definition of logarithm
:
(t / 5750) = log base (1/2) 0.20 = 2.3219
:
t = 5750 * 2.3219 = 13350.925
:
*******************************************
The object is 13,350 years old
*******************************************
: